A few years ago, a total of 2687 thousand people lived in las vegas, and sacramento. Sacramento had 285 thousand more residents than Las vegas. what was the the population in both las vegas and sacramento?
Answer by CubeyThePenguin(3113)

S = Sacramento's population
L = Las Vegas' population

S + L = 2,687,000
S = L + 285,000

Plug in the second equation into the first.

2L + 285,000 = 2,687,000

There are 1,201,000 people in Las Vegas and 1,486,000 people in Sacramento.